John is three times as old as Mary and the sum of their ages is 24. find their ages
Answer by stanbon(75887)

John is three times as old as Mary and the sum of their ages is 24. find their ages
---
Equations:
j = 3*m
j+m = 24
----
xubstitute for "j" and solve for "m":
3m + m = 24
4m= 24
Mary = 6 years
----
solve for "j":
j = 3m
John = 3*6
John = 18 years
====================
cheers,
Stan H.